# Learning Encoding-Decoding Direction Pairs to Unveil Concepts of Influence in Deep Vision Networks

### v2

@@ 7.4.3 Evaluation of the Encoding Directions: Influence
We use RCAV RCAV to assess the ability of our method to identify concepts influential to model predictions and relate concept influence to the network’s concept sensitivity. In the ablation studies, two metrics summarize the results: Significant Direction Count (SDC) and Significant Class-Direction Pairs (SCDP). SDC represents the number of signal vectors that significantly influence at least one model class, while SCDP tallies class-direction pairs where the signal vectors significantly affect the class. Let $S_{C,i,k}$ denote the sensitivity of the model to concept $i$ when predicting class $k$. When comparing against unsupervised baselines, we use the following influence metric $\mathcal{I}^{1}$ which is defined as:
@@ 7.5 Interpretability and Influence Comparison with Established Unsupervised Baselines
The previous unsupervised state-of-the-art (PCA; SVD; NMFCAV; CRAFT) has demonstrated effectiveness in revealing interpretable directions in the latent space of deep networks. However, as outlined in Section A.12, these approaches employed a class-conditioned protocol, tying their analyses to specific class labels. In this work, we instead operate under a class-agnostic protocol, which seeks concepts across the entire latent space, independent of class. This fundamental difference in the approaches renders a direct performance comparison with previous implementations inapplicable. Therefore, we utilize PCA and NMF, the core matrix decomposition methods upon which these prior works were built, as our principal baselines for a fair, classical comparison. We also move beyond subjective human ratings typically used for evaluation, and instead develop a rigorous, best-effort quantitative framework to compare interpretability and influence with these baselines. In Section A.15, we also compare with the methods we extend (UIBE; CBE), but under a different protocol for a fair comparison. For each network, we consider three sizes for the clustering, namely $I=3/4D$, $I=7/8D$, and $I=D$. When $I=D$, NMF trivially resolves to the natural latent space basis, i.e., the neuron directions. When studying EfficientNet, we disregard NMF since this network uses the SiLU activation function (silu), which allows negative feature activations, rendering NMF non-applicable. Tables 4,5, 6, 7, and 8 summarize the comparative analysis for ResNet18, EfficientNet, Inception-v3, VGG16, and ResNet50, respectively. In the following subsections, we have a detailed discussion of those tables.
7.5.1 Analysis for ResNet18 trained on Places365
@@
7.5.2 Analysis for EfficientNet trained on ImageNet
@@
7.5.3 Analysis for Inception-v3 trained on ImageNet
@@
7.5.4 Analysis for VGG16 trained on ImageNet
@@
Interpretability: Unlike the rest of the architectures, in classification metrics, NMF and the EDDP variants perform on par. For example, when $I=448$, NMF attains a Precision score of $0.62$ exceeding EDDP-U’s $0.59$. In Recall and F1 score terms, EDDP-U achieves the highest scores of $0.18$ and $0.26$, respectively, while for NMF these numbers are $0.16$ and $0.25$. In terms of AP, NMF scores $0.31$ while both EDDP variants attain a score of $0.30$. In most cases, all of NMF, the natural basis, and EDDP substantially surpass the clustering of PCA in the classification metrics.
Even though EDDP, NMF and the natural latent space basis score similarly in Precision terms, when monosemanticity is measured via the CLIP embedding distances, the EDDP variants achieve the best scores, often by a substantial margin. For instance, when $I=384$, NMF achieves an $\mathcal{M}$ score of $7.44$, PCA a score of $7.85$, while EDDP-U a score of $6.88$.
@@
Influence: Regarding the sensitivity of VGG16 with respect to the concepts, the network is the most sensitive towards the directions learned with EDDP-U ($\mathcal{I}^{1}$ score of $0.91$), followed by EDDP-C ($\mathcal{I}^{1}$ score of $0.9$) and last, by NMF, the natural basis, or PCA with $\mathcal{I}^{1}$ scores close to $0.87$.
7.5.5 Analysis for ResNet50 trained on Moments in Time
@@
Clustering Quality: Once again, in clustering quality terms, the EDDP variants demonstrate both larger Coverage and lower Redundancy compared to the rest of the approaches, often with a large margin. When $I=1792$, the coverage score for EDDP-C is $0.88$ while for NMF it is $0.42$. The gap between the two is more than twice the score of NMF. Similarly, the redundancy score of EDDP-C is $2.62$ whereas NMF’s is $6.02$, with a similar conclusion regarding the gap between the two.
@@
Influence: In terms of the concept sensitivity metric $\mathcal{I}^{1}$, NMF, EDDP-C, EDDP-U and the natural basis perform on par, with an approximate score of $0.85$. On the contrary, PCA achieves a score of $0.88$, which brings it further ahead of the competition.
@@ 7.5.6 Discussion
From the previous experiments, we can draw the following, mostly straightforward, conclusions: a) EDDP’s directional clusters are monosemantic, often surpassing the unsupervised baselines by a significant margin. b) EDDP also excels in terms of F1 score, where EDDP-U is consistently ranked first among the competition, and EDDP-C is often ranked second. c) In terms of segmentation and specifically $\mathcal{S}^{1}$, EDDP-U is consistently ranked first, except for EfficientNet in which the clustering of PCA scores better, yet with substantially inferior performance in most of the other interpretability metrics. In the same metric, EDDP-C is often ranked second. d) In terms of discovering a diverse set of concepts ($\mathcal{S}^{2}$) the EDDP variants score the best, often by a large margin, except for one case in VGG16 where they score on par with the best method. e) In terms of influence, in all cases except VGG16, the EDDP variants are ranked last. However, in many cases, the respective score is not far behind the competition.
In most cases, and in a broader sense, we observe that interpretability is somewhat negatively correlated with influence. Less interpretable directions may represent complex interactions (i.e. correspond to less pure human-understandable concepts) and, as a result, they are more impactful on the network predictions, in the sense that they affect many output classes. On the contrary, more interpretable directions may have a more targeted impact on specific output classes (sae-scaling), and thus score lower in the respective metric.
@@ 7.6.1 Interpretability Losses
The motivation behind the loss terms that we introduced in Section 4.2 is to sustain or improve the interpretability of the clustering when moving from an orthogonal decoding direction set to a non-orthogonal one. In Table 9 we present interpretability metrics for the cases when we learn the directions with $\mathcal{L}^{fs}$, both $\mathcal{L}^{fs}$ and $\mathcal{L}^{eac}$, or without any of them. We also provide metrics for the case of orthogonal decoding directions (marked as Ortho). The latter is equivalent to learning with the method of UIBE using the Augmented Lagrangian loss of Section 4.5. We provide results regarding the directions learned for ResNet18 and EfficientNet. In these experiments, we exclude any form of Uncertainty Region Alignment.
@@ 7.7 Insights on Interpretability and Influence in comparison to Supervised Direction Learning
Table 12: Comparison of EDDP (with CFM) with supervised direction learning IBD. Comparing between: a) standard EDDP, b) combined concept detectors (Linear-OR) c) EDDP with the thresholds of concept detectors learned with supervision in a post initial direction learning step (/w sup thres), and d) IBD: a set of classifiers learned in a supervised way. The network here is ResNet18 and $I=512$.
@@
In this section, we compare our method with supervised direction learning, aiming to draw insights regarding interpretability and influence of concept directions. For this reason, we experiment with the last convolutional layer of ResNet18 trained on Places365. We consider our method with the constrained feature manipulation strategy, and two supervised learning approaches: a) IBD (IBD) for supervised learning of decoding directions (i.e. concept detectors) and b) Pattern-CAVs (PCAVs) (PCAV) for supervised estimation of concept encoding directions. The supervised concept detectors and PCAVs are learned for the labels assigned to EDDP’s concept detectors at the direction labeling phase (Section 3.4).
@@
Influence: In the concept influence analysis, we aim to answer whether PCAVs, that correspond to directions of high interpretability due to learning them with supervision, are more influential to the network predictions than signal vectors. Since Network Dissection can assign identical labels to multiple Encoding-Decoding Direction Pairs, direct comparison with Pattern-CAVs becomes less straightforward. We propose the following metric to help us draw some conclusions. Let $j\in\{0,1,...,N_{l}-1\}$ index EDDP’s signal vectors sharing the same concept label $l$, and $S^{l}_{j,k}$ represent the RCAV sensitivity score of the network relative to the $j$-th signal vector of label $l$ when predicting class $k$. Similarly, let $S^{l}_{P,k}$ denote the sensitivity score of the network relative to the Pattern-CAV for the same label and class. Inspired by RCAV, we regard signal vectors as noise vectors and compare the sensitivity of the network with respect to PCAVs and signal vectors that share the same label. We define a metric $\mathcal{I}^{2}$, which when above 0.5 indicates that PCAVs have more influence than signal vectors on the network predictions at the statistical significance level of $\theta=0.05$ with Bonferroni correction:
@@
For RCAV $\alpha=5.0$, $\mathcal{I}^{2}$ is equal to $0.34$, which is less than $0.5$ indicating the Pattern-CAVs are less influential on the network predictions than signal vectors. This outcome could also be linked to the discussion of Section 7.5.6. It looks like less interpretable directions (in that case EDDP’s) have more influence on the network predictions compared to more interpretable ones.
@@ 7.11.4 Signal Vector Faithfulness to the Watermark Concept
Leveraging our knowledge of which images contain the watermark, we can determine the watermark’s encoding direction by calculating the watermark’s Pattern-CAV (PCAV). Subsequently, we quantify the directional alignment between the Pattern-CAV, which was learned with supervision, and the signal vector, which was learned without. We found that the cosine similarity between these two is $0.99$, indicating estimations that agree.
@@ 8 Limitations
Our method builds on the linear representation hypothesis, assuming that concepts are encoded as directions in the latent space. While this view aligns with several prior works, alternative formulations such as multidimensional concept discovery (mcd) or linear subspaces (RelevantSubspaces) exist but fall outside the scope of our study. Additionally, we constrain our experiments to settings where the number of discovered concepts is smaller than the embedding dimensionality. As such, we do not compare with methods like Sparse Autoencoders (SAEs) (SAE1; SAE2; SAE3) that assume an overcomplete representation. Even though we restricted our experiments as such, future studies may consider this analysis, since our method is not technically limited in this regard. While we evaluated our approach across multiple CNN architectures, we did not investigate applications to Vision Transformers (ViTs) (dino; mae), despite existing literature suggesting that similar linear assumptions may apply in that context (xai-directions)—this remains a promising direction for future work. Moreover, unlike some methods from unsupervised state-of-the-art that operate on class-specific image subsets constructed from model predictions, we applied our method to unlabeled data without class filtering, due to the lack of dense annotations in the training sets. Nevertheless, our framework could, in principle, be adapted for class-conditional analysis. Finally, our study aimed to minimize hyperparameter tuning by relying on prior work (UIBE; CBE) and empirical preliminary experiments; however, further tuning or exploration of additional hyperparameters may yield improvements in clustering quality, interpretability, or influence (Section A.13).
@@ Appendix A Appendix
  • A.10: Details on the Faithfulness Assessment of Encoding-Decoding Direction Pairs
  • A.11: Additional Experiments on the Faithfulness Assessment of the Encoding-Decoding Direction Pairs
  • A.12: Details for the Experiments on Deep Image Classifiers
  • A.13: Hyper-Parameter Study with Respect to Target Separation Margin
  • A.14: Ablation Study with respect to Filter-Signal Orthogonality Loss
  • A.15: Comparison with Unsupervised Interpretable Basis Extraction and Concept-Basis Extraction in Practical Experiments
  • A.16: More Qualitative Segmentations and Statistics for Evaluating the Interpretability of the Concept Detectors
  • A.17 More Global Model Explanations via Concept Sensitivity Testing
  • A.18 More Local Explanations with Concept Contribution Maps
  • A.19 Details on Counterfactual Explanation Example
  • A.20 Details on Toy Model Correction
  • A.21 Details on Computational Resources
  • A.22 Relation to Sparse AutoEncoders (SAEs)

### v3

@@ 7.4.3 Evaluation of the Encoding Directions: Influence
We use RCAV Pfau et al. (2020) to assess the ability of our method to identify concepts influential to model predictions and relate concept influence to the network’s concept sensitivity. In the ablation studies, two metrics summarize the results: Significant Direction Count (SDC) and Significant Class-Direction Pairs (SCDP). SDC represents the number of signal vectors that significantly influence at least one model class, while SCDP tallies class-direction pairs where the signal vectors significantly affect the class. Let $S_{C,i,k}$ denote the sensitivity of the model to concept $i$ when predicting class $k$. When comparing against unsupervised baselines, we use the following influence metric $\mathcal{I}^{1}$ which is defined as:
@@ 7.5 Interpretability and Influence Comparison with Established Unsupervised Baselines
The previous unsupervised state-of-the-art (Graziani et al., 2023a; b; Zhang et al., 2021; Fel et al., 2023b) has demonstrated effectiveness in revealing interpretable directions in the latent space of deep networks. However, as outlined in Section A.13, these approaches employed a class-conditioned protocol, tying their analyses to specific class labels. In this work, we instead operate under a class-agnostic protocol, which seeks concepts across the entire latent space, independent of class. This fundamental difference in the approaches renders a direct performance comparison with previous implementations inapplicable. Therefore, we utilize PCA and NMF, the core matrix decomposition methods upon which these prior works were built, as our principal baselines for a fair, classical comparison. We also move beyond subjective human ratings typically used for evaluation and instead develop a rigorous, best-effort quantitative framework to compare interpretability and influence with these baselines, with a more detailed discussion on how our pipeline captures human judgment provided in Section A.10.
In Section A.16, we also compare with the methods we extend (Doumanoglou et al., 2023; 2024), but under a different protocol for a fair comparison. For each network, we consider three sizes for the clustering, namely $I=3/4D$, $I=7/8D$, and $I=D$. It is worth noting that when $I=D$, NMF trivially resolves to the natural latent space basis, i.e., the neuron directions. Additionally, when studying EfficientNet, we disregard NMF since this network uses the SiLU activation function (Ramachandran et al., 2017), which allows negative feature activations, rendering NMF non-applicable. Tables 4, 5, 6, 7, and 8 summarize the comparative analysis for ResNet18, EfficientNet, Inception-v3, VGG16, and ResNet50, respectively. In the following subsections, we provide a detailed discussion of these tables.
7.5.1 Case Study: ResNet18 / Places365
@@
7.5.2 Case Study: EfficientNet / ImageNet
@@
7.5.3 Case Study: Inception-v3 / ImageNet
@@
7.5.4 Case Study: VGG16 / ImageNet
@@
Interpretability: Unlike the other architectures, in classification metrics, NMF and the EDDP variants perform on par. For example, when $I=448$, NMF attains a Precision score of $0.62$ exceeding EDDP-U’s $0.59$. In Recall and F1 score terms, EDDP-U achieves the highest scores of $0.18$ and $0.26$, respectively, while for NMF these numbers are $0.16$ and $0.25$. In terms of AP, NMF scores $0.31$ while both EDDP variants attain a score of $0.30$. In most cases, all of NMF, the natural basis, and EDDP substantially surpass the clustering of PCA in the classification metrics.
Even though EDDP, NMF, and the natural latent space basis score similarly in terms of Precision, if monosemanticity is measured via the CLIP embedding distances, the EDDP variants achieve the best scores, often by a substantial margin. For instance, when $I=384$, NMF achieves an $\mathcal{M}$ score of $7.44$, PCA a score of $7.85$, while EDDP-U a score of $6.88$.
@@
Influence: Regarding the sensitivity of VGG16 with respect to the concepts, the network is most sensitive to the directions learned with EDDP-U ($\mathcal{I}^{1}$ score of $0.91$), followed by EDDP-C ($\mathcal{I}^{1}$ score of $0.9$), and finally, by NMF, the natural basis, or PCA, with $\mathcal{I}^{1}$ scores close to $0.87$.
7.5.5 Case Study: ResNet50 / Moments in Time
@@
Clustering Quality: Once again, in terms of clustering quality, the EDDP variants demonstrate both larger Coverage and lower Redundancy compared to the rest of the approaches, often with a large margin. When $I=1792$, the coverage score for EDDP-C is $0.88$ while for NMF it is $0.42$. The gap between the two is more than twice the score of NMF. Similarly, the redundancy score of EDDP-C is $2.62$ whereas NMF’s is $6.02$, with a similar conclusion regarding the gap between the two.
@@
Influence: In terms of the concept sensitivity metric $\mathcal{I}^{1}$, NMF, EDDP-C, EDDP-U and the natural basis perform on par, with an approximate score of $0.85$. On the contrary, PCA achieves a score of $0.88$, which puts it further ahead of the competition.
@@ 7.5.6 Discussion
From the previous experiments, we can draw the following, mostly straightforward conclusions: a) EDDP’s directional clusters are monosemantic, often surpassing the unsupervised baselines by a significant margin. b) EDDP also excels in terms of F1 score, where EDDP-U is consistently ranked first among the competition, and EDDP-C is often ranked second. c) In terms of segmentation and specifically $\mathcal{S}^{1}$, EDDP-U is consistently ranked first, except for EfficientNet, in which the clustering of PCA scores better, yet with substantially inferior performance in most of the other interpretability metrics. In the same metric, EDDP-C is often ranked second. d) In terms of discovering a diverse set of concepts ($\mathcal{S}^{2}$) the EDDP variants score the best, often by a large margin, except for one case in VGG16 where they score on par with the best method. e) In terms of influence, in all cases except VGG16, the EDDP variants are ranked last. However, in many cases, the respective score is not far behind the competition.
In most cases, and in a broader sense, we observe that interpretability is somewhat negatively correlated with influence. Less interpretable directions may represent complex interactions (i.e. correspond to less pure human-understandable concepts) and, as a result, they are more impactful on the network predictions, in the sense that they affect many output classes. On the contrary, more interpretable directions may have a more targeted impact on specific output classes (Gao et al., 2025), and thus score lower in the respective metric.
@@ 7.6.1 Interpretability Losses
The motivation behind the loss terms that we introduced in Section 4.2 is to sustain or improve the interpretability of the clustering when moving from an orthogonal decoding direction set to a non-orthogonal one. In Table 9 we present interpretability metrics for the cases when we learn the directions with $\mathcal{L}^{fs}$, both $\mathcal{L}^{fs}$ and $\mathcal{L}^{eac}$, or without any of them. We also provide metrics for the case of orthogonal decoding directions (marked as Ortho). The latter is equivalent to learning with the method of Doumanoglou et al. (2023) using the Augmented Lagrangian loss of Section 4.5. We provide results regarding the directions learned for ResNet18 and EfficientNet. In these experiments, we exclude any form of Uncertainty Region Alignment.
@@ 7.7 Insights on Interpretability and Influence in comparison to Supervised Direction Learning
Table 12: Comparison of EDDP (with CFM) with supervised direction learning Zhou et al. (2018). Comparing between: a) standard EDDP, b) combined concept detectors (Linear-OR) c) EDDP with the thresholds of concept detectors learned with supervision in a post initial direction learning step (/w sup thres), and d) IBD: a set of classifiers learned in a supervised way. The network here is ResNet18 and $I=512$.
@@
In this section, we compare our method with supervised direction learning, aiming to draw insights regarding interpretability and influence of concept directions. For this reason, we experiment with the last convolutional layer of ResNet18 trained on Places365. We consider our method with the constrained feature manipulation strategy, and two supervised learning approaches: a) IBD (Zhou et al., 2018) for supervised learning of decoding directions (i.e. concept detectors) and b) Pattern-CAVs (PCAVs) (Pahde et al., 2024) for supervised estimation of concept encoding directions. The supervised concept detectors and PCAVs are learned for the labels assigned to EDDP’s concept detectors at the direction labeling phase (Section 3.4).
@@
Influence: In the concept influence analysis, we aim to answer whether PCAVs, that correspond to directions of high interpretability due to learning them with supervision, are more influential to the network predictions than signal vectors. Since Network Dissection can assign identical labels to multiple Encoding-Decoding Direction Pairs, direct comparison with Pattern-CAVs becomes less straightforward. We propose the following metric to help us draw some conclusions. Let $j\in\{0,1,...,N_{l}-1\}$ index EDDP’s signal vectors sharing the same concept label $l$, and $S^{l}_{j,k}$ represent the RCAV sensitivity score of the network relative to the $j$-th signal vector of label $l$ when predicting class $k$. Similarly, let $S^{l}_{P,k}$ denote the sensitivity score of the network relative to the Pattern-CAV for the same label and class. Inspired by RCAV, we regard signal vectors as noise vectors and compare the sensitivity of the network with respect to PCAVs and signal vectors that share the same label. We define a metric $\mathcal{I}^{2}$, which, when above 0.5, indicates that PCAVs have more influence than signal vectors on the network predictions at the statistical significance level of $\theta=0.05$ with Bonferroni correction:
@@
For RCAV $\alpha=5.0$, $\mathcal{I}^{2}$ is equal to $0.34$, which is less than $0.5$, indicating that the Pattern-CAVs are less influential on the network predictions than signal vectors. This outcome could also be linked to the discussion in Section 7.5.6. It appears that less interpretable directions (in this case, EDDP’s) have more influence on the network predictions compared to more interpretable ones.
@@ 7.11.4 Signal Vector Faithfulness to the Watermark Concept
Leveraging our knowledge of which images contain the watermark, we can determine the watermark’s encoding direction by calculating the watermark’s Pattern-CAV (Pahde et al., 2024). Subsequently, we quantify the directional alignment between the Pattern-CAV, which was learned with supervision, and the signal vector, which was learned without. We found that the cosine similarity between these two is $0.99$, indicating estimations that agree.
@@ 8 Limitations
Our method builds on the linear representation hypothesis, assuming that concepts are encoded as directions in the latent space. While this view aligns with several prior works, alternative formulations such as multidimensional concept discovery (Vielhaben et al., 2023) or linear subspaces (Chormai et al., 2024) exist but fall outside the scope of our study. Additionally, we constrain our experiments to settings where the number of discovered concepts is smaller than the embedding dimensionality. As such, we do not compare with methods like Sparse Autoencoders (SAEs) (Lim et al., 2024; Cunningham et al., 2024; Sharkey et al., 2022) that assume an overcomplete representation. Even though we restricted our experiments as such, future studies may consider this analysis, since our method is not technically limited in this regard. While we evaluated our approach across multiple CNN architectures, we did not investigate applications to Vision Transformers (ViTs) (Caron et al., 2021; He et al., 2022), despite existing literature suggesting that similar linear assumptions may apply in that context (Kaltampanidis et al., 2025)—this remains a promising direction for future work. Moreover, unlike some methods from unsupervised state-of-the-art that operate on class-specific image subsets constructed from model predictions, we applied our method to unlabeled data without class filtering, due to the lack of dense annotations in the training sets. Nevertheless, our framework could, in principle, be adapted for class-conditional analysis. Finally, our study aimed to minimize hyperparameter tuning by relying on prior work (Doumanoglou et al., 2023; 2024) and empirical preliminary experiments; however, further tuning or exploration of additional hyperparameters may yield improvements in clustering quality, interpretability, or influence (Section A.14).
@@ 9 Conclusion
Acknowledgments
This work has been supported by the EC funded Horizon Europe Framework Programme: CAVAA Grant Agreement 101071178.
@@ Appendix A Appendix
  • A.10: Approaching Human Evaluations
  • A.11: Details on the Faithfulness Assessment of Encoding-Decoding Direction Pairs
  • A.12: Additional Experiments on the Faithfulness Assessment of the Encoding-Decoding Direction Pairs
  • A.13: Details for the Experiments on Deep Image Classifiers
  • A.14: Hyper-Parameter Study with Respect to Target Separation Margin
  • A.15: Ablation Study with respect to Filter-Signal Orthogonality Loss
  • A.16: Comparison with Unsupervised Interpretable Basis Extraction and Concept-Basis Extraction in Practical Experiments
  • A.17: More Qualitative Segmentations and Statistics for Evaluating the Interpretability of the Concept Detectors
  • A.18 More Global Model Explanations via Concept Sensitivity Testing
  • A.19 More Local Explanations with Concept Contribution Maps
  • A.20 Details on Counterfactual Explanation Example
  • A.21 Details on Toy Model Correction
  • A.22 Details on Computational Resources
  • A.23 Relation to Sparse AutoEncoders (SAEs)